I have a 7 layer torte that I am making. Each layer is 8 inches in diameter with cream topping on each layer. The recipe I have for the cream topping will cover 55 inches of torte. How many batches of cream topping will I have to make?
Use n = 3.14 approximate value.
I have no idea how to begin this equation. Thank you for your assistance!!

Answer by rfer(16322)

find the area of each and then the total area
4^2*3.14=50.24 sq in
50.24*7=351.68
351.68/55=6.39 batches
It is necessary to make 7 batches